# Scalable Training of Trustworthy and Energy-Efficient Predictive Graph Foundation Models for Atomistic Materials Modeling: A Case Study with HydraGNN

### v3

@@ Abstract
We present our work on developing and training scalable, trustworthy, and energy-efficient predictive graph foundation models (GFMs) using HydraGNN, a multi-headed graph convolutional neural network architecture. HydraGNN expands the boundaries of graph neural network (GNN) computations in both training scale and data diversity. It abstracts over message passing algorithms, allowing both reproduction of and comparison across algorithmic innovations that define nearest-neighbor convolution in GNNs. This work discusses a series of optimizations that have allowed scaling up the GFMs training to tens of thousands of GPUs on datasets that consist of hundreds of millions of graphs. Our GFMs use multi-task learning (MTL) to simultaneously learn graph-level and node-level properties of atomistic structures, such as energy and atomic forces. Using over 154 million atomistic structures for training, we illustrate the performance of our approach along with the lessons learned on two state-of-the-art United States Department of Energy (US-DOE) supercomputers, namely the Perlmutter petascale system at the National Energy Research Scientific Computing Center and the Frontier exascale system at Oak Ridge Leadership Computing Facility. The HydraGNN architecture enables the GFM to achieve near-linear strong scaling performance using more than 2,000 GPUs on Perlmutter and 16,000 GPUs on Frontier. Hyperparameter optimization (HPO) was performed on over 64,000 Graphic Compute Dies (GCDs) on Frontier to select GFM architectures with high accuracy. Early stopping was applied on each GFM architecture for energy awareness in performing such an extreme-scale task. Each HPO trial was ranked based on both accuracy and energy consumption. The training of an ensemble of highest-ranked GFM architectures (selected with judicious balance between accuracy and energy consumption) continued until convergence to establish uncertainty quantification (UQ) capabilities with ensemble learning. Our contributions establish core capabilities for rapidly developing, training, and deploying further GFMs using large-scale computational resources to enable AI-accelerated materials discovery and design in a trustworthy and energy-efficient manner.
Index Terms:
Machine Learning, Distributed Data Parallelism, Graph Neural Networks, Message Passing, Graph Foundation Models, Large-Scale Data Processing for Machine Learning, Atomistic Materials Modeling
@@ I Introduction
Discovery of new materials with desired properties, accurate predictions of a material’s behavior throughout its entire lifespan is crucial to fundamental scientific progress in energy generation, transportation, electronics, and information technology [1]. Machine learning (ML) has shown great potential in accelerating the screening and pre-selection of materials for further experimental testing. In particular, deep learning (DL) models have shown the ability to effectively capture relevant underlying relationships due to the arrangement of atoms of different constituents within various atomistic structures [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. DL models can be trained on data generated from experiments and/or first-principles calculations and then used to predict the properties of interest to be used as new inputs. The inference takes only a fraction of the time it would take to run an experiment or a full first-principles calculation while still producing sufficiently accurate results. This drastic reduction in time to predict material properties using atomistic information results in a promising path towards accelerating material discovery and design [13, 14].
@@
However, generating vast volumes of experimental and/or first-principles training data is impractical even with sophisticated experimental facilities and powerful supercomputers. Recently, foundation models (FMs) have demonstrated the capability to navigate around the challenge: once pre-trained over a large volume of available open-source data [15], an FM holds the promise to provide a jump-start to refined models by fine-tuning on significantly smaller amounts of data for customized applications (also called downstream tasks). Reducing the number of simulations and/or experiments for generating domain-specific training data also drastically reduces the energy costs of developing domain-specific DL models.
@@
While state-of-the-art language-based FMs with a transformer architecture have reached promising results in several domains [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], they fail to capture important topological aspects of the atomistic structures. Therefore, alternative DL architectures that are better suited to retain important topological aspects at the atomistic scale need to be considered for the development of trustworthy (interpreted as simultaneously accurate and highly confident) FMs for materials using atomistic scale information.
@@
Since atomistic material structures for a generic type of compound can be mapped onto a graph (where atoms can be treated as nodes and interatomic bonds as edges), graph foundation models (GFMs) [32, 33], which are FMs that operate on data structures as graphs, are the candidate of choice for these applications. Currently, GFMs proposed in the literature are developed by training graph neural network (GNN) architectures on a sufficiently large and comprehensive dataset for the domain of interest. While a few efforts have already been undertaken to develop GFMs for atomistic materials modeling applications [34, 35, 36, 37], the existing work is still at an incipient stage. Current efforts do not yet ensure that their proposed approach achieves trustworthiness. Moreover, serious concerns have been recently raised about the unaffordable amount of energy requested by proliferating AI centers due to computationally intensive tasks of training large FMs on large volumes of data. Therefore, the development of energy-efficient methodologies for a sustainable development of FMs is of paramount importance.
@@
The work described in this manuscript developed scalable and trustworthy supervised GFMs for the simultaneous prediction of energies and atomic forces with careful attention to consider energy consumption. The GFMs have been constructed using HydraGNN [38, 39], a fully scalable GNN architecture developed at ORNL. Experiments were conducted on two large US-DOE supercomputers: the Perlmutter petascale machine at National Energy Research Scientific Computing Center (NERSC) and the Frontier exascale system at Oak Ridge Leadership Computing Facility (OLCF).
@@
The remaining content of the paper is organized as follows. We discuss the current state of the art and introduce HydraGNN in Section II. In Section III, we discuss our approach toward developing a scalable framework and list the different optimizations for scalable training. We discuss our use of large-scale HPO to develop a trustworthy GFM. Section IV shows the performance of different components of this work: reading large data, scaling the training process, performing HPO at large scale, and measuring epistemic uncertainty using an ensemble of HPO trials with high predictive accuracy and low energy costs. We conclude our study and discuss future work in Section V.
II Current State of the Art
II-A GNN training open-source atomistic data
@@
To date, there have been a few approaches proposed in the literature to develop GFMs for atomistic materials modeling. In [34], the authors proposed a multi-modal approach where multiple encoding DL architectures are trained on different types of data representations and describing different physical quantities. The models are aligned to each other through a penalization term in the training loss function that forces latent vectors from each embedding space to coincide. However, the datasets used comprise only organic molecules, which allows them to cover only a relatively small set of natural elements on the periodic table.
@@
In [42], the authors propose a new kernel function for GNNs for modeling potential energy surfaces, which exploits chemical species information, and illustrate its efficacy in the context of transferable GFMs by pretraining the GFM on the Open Catalyst 2020 (OC2020) dataset [43] and illustrating its performance on a set of fine-tuning tasks. Since the GFM was pretrained only on the OC2020 dataset, the applicability of this GFM is restricted to inorganic compounds.
While not explicitly presented by their developers as GFMs, there have been other models that cover broader sets of elements of the periodic table compared to the approaches mentioned in the previous paragraph. In [35], the authors built a GNN model using MTL for simultaneous predictions of several material properties by training the GNN model on multiple datasets, including OC2020 [43] and Open Catalyst 2022 (OC2022) [44]. However, the approach considers only a single GNN architecture without performing HPO. Moreover, the set of parameters in the GNN model is relatively small, in the order of few millions of parameters, which limits the attainable accuracy on large volumes of data.
In [37], the authors studied the scaling behavior of 2D molecular GNNs under varied settings of depth, width, number of molecules, number of labels, the diversity in dataset, and architectural choice. The authors showed that supervised pretraining of large GNNs on molecular datasets provides a rich fingerprint embedding, which is useful for 38 downstream tasks. Even if this work very systematically studied the effect of GNN model size over the predictive performance in the pretraining and fine-tuning stage with many and diverse downstream tasks, the work has two important limitations: it only considers 2D graphs and it addressed only organic compounds.
Several UQ methods have been applied to GNNs [45], including Bayesian GNNs [46], prediction interval methods [47], and deep ensemble methods [48]. Bayesian methods are theoretically rigorous but challenging to scale to high-dimensional data. Prediction interval methods are cost-effective but often require tedious tuning of heuristic parameters. For UQ, we use deep ensemble methods as a compromise between cost and performance to quantify uncertainty in our GFMs.
Compared to the scientific contributions mentioned above, our work distinguishes itself by leveraging extreme scale supercomputing resources to ensure trustworthiness of the GFMs by performing (i) a systematic large scale HPO across a broad set of GNN architectures and (ii) a large scale ensemble learning (EL) for UQ.
II-B Scalability and GPU optimization for GNN training
@@
These are useful conclusions for optimization of GNN training on large graphs (i.e., with millions of nodes), but need to be re-evaluated for our datasets. Training on large graphs can be highly sensitive to the splitting scheme used to partition the graph into subgraphs and to distribute them among processors. For atomistic materials modeling applications addressed in our work, the graph samples are small (with at most a few hundreds of nodes). For GNN convolutions in particular, convolution on a batch of samples will have a much more local, block diagonal structure. Therefore, throughput should be less sensitive to the choice of atomistic structures per batch.
@@
II-C HydraGNN
The complexity of the physics and the scale at which atomistic structures must be studied in response to US-DOE needs in materials science makes it compelling to develop GNN capabilities that simultaneously satisfy several important algorithmic and computer science requirements. To effectively respond to the scientific needs of the US-DOE, a GNN architecture must provide: (1) capabilities to read and process data from multiple sources simultaneously, (2) flexibility to support diverse DOE-relevant scientific applications, (3) capabilities to scale the training on leadership class supercomputing facilities, (4) portability across heterogeneous computing environments, (5) continuous software maintenance by ensuring support and compatibility with upgraded software dependencies, (6) maintained documentation to support new users across a broad set of international institutions.
While several GNN architectures have been made available as open-source tools to the scientific community in the last few years [55, 56, 57, 58], none of these tools completely satisfies the above requirements. Moreover, including missing capabilities on these well-established GNN libraries requires invasive and laborious modifications for software re-design. These challenges arising from existing GNN implementations motivated our effort in developing HydraGNN [38, 39]. In response to the US-DOE scientific needs, HydraGNN provides:
  • multi-task learning (MTL) capabilities to process multi-source, multi-fidelity data [59]
  • object-oriented programming capabilities to use different MPNN layers [60], which allows flexible switching between different message policies based on the scientific needs of the specific application at hand, as well treating the MPNN layer as a tunable categorical hyperparameter with HPO
  • invariant and equivariant features that reduce computational redundancy and time-to-solution [61], therefore contributing to energy saving
  • scalable input/output (I/O) data management techniques to efficiently scale the training of GNN models on millions of data samples using thousands of GPUs on supercomputing facilities [62]
  • portable capabilities that allow conveniently running the GNN training on diverse computing platforms with different hardware and software specifications
@@ II-C HydraGNN
III Our Contribution
The work described in this manuscript is the first-of-its- kind large-scale training of GFMs for atomistic materials modeling using over 154 million atomistic structures as data samples and using over $91\%$ of the exascale supercomputer Frontier. We have employed three key techniques for developing a scalable and trustworthy GFM: 1) scalable data management using a scientific data management library and an in-memory data store, 2) scalable HPO that uses asynchronous Bayesian optimization for efficiently managing computing resources, and 3) ensemble methods for UQ that allows model generalization and concurrently training multiple models. These three advancements collectively enhance the robustness, efficiency, and scalability of the GNN training process.
@@ III-A Data Aggregation
TABLE I: Overview of Datasets used for training HydraGNN
Using large datasets for GFM training can enhance generalizability and ensure resilience to data variance issues that typically arise during downstream tasks. To this end, we aggregated five open-source atomistic materials modeling datasets that are extremely diverse in terms of chemical composition, atomistic configurations, and number of atoms in the system. These datasets, as listed in Table I, are: ANI1x, QM7x, OC2020, OC2022, and MPTrj.
  • ANI1x [67] consists of over 4,956,005 conformations derived from up to 57 thousand distinct molecular configurations containing the C, H, N, and O chemical elements
  • QM7x [68] is a comprehensive dataset of 42 physicochemical properties for approximately 4.2 million equilibrium and non-equilibrium structures of small organic molecules with up to seven non-hydrogen atoms from the C, N, O, S, Cl chemical elements
@@
Each dataset is unique for the chemical compositions and the number of atoms in the atomistic structures of the compounds described. Figure 1 shows the distribution of the number of atoms and graph edges per atomistic structure for each dataset. For the MPTrj dataset, approximately half of the atomistic structures are relatively small in size. On the other hand, the OC2020 and OC2022 datasets consist of a more even distribution of atomistic structures with different sizes and edge counts, with the larger structures consisting of over 400 atoms and over 12,500 edges. In total, the data used for training, validating, and testing our GFM consisted of over 154 million atomistic structures that consume 5.3 Terabytes of storage space. These datasets were pre-processed using a scientific data management library into a common format for efficient storage and I/O, as discussed in Section III-C.
@@
Figure 2 provides the heatmap that illustrates the frequency of occurrence of each element of the periodic table across the entire dataset that results from the aggregation of the datasets ANI1x, QM7-X, OC2020, OC2022, and MPTrj. We notice that there is an under-representation of the elements from the transition metal groups. This is partially due to the fact that first-principles calculations for materials including these elements (e.g., alloys) are more computationally expensive to run, and thus are available in smaller volumes with respect to first-principles data for other classes of materials.
@@
III-B Data Cleaning and Pre-processing
Some of the atomistic structures were determined to have unrealistic values for atomic forces (on the order of 20,000 eV/angstrom). These probably corresponded to configurations visited at early stages of energy minimization during the generation of these datasets. To eliminate these outliers, we first applied a data cleaning operation in which we discarded all atomistic structures with an $L2$-norm (also known as spectral norm) of the force tensor above 100 eV/angstrom to ensure that these data samples did not affect the training of our GFMs. The number of data samples removed from each dataset by this filtering operation is reported in Table II.

### v4

@@ Abstract
We present our work on developing and training scalable, trustworthy, and energy-efficient predictive graph foundation models (GFMs) using HydraGNN, a multi-headed graph convolutional neural network architecture. HydraGNN expands the boundaries of graph neural network (GNN) computations in both training scale and data diversity. It abstracts over message passing algorithms, allowing both reproduction of and comparison across algorithmic innovations that define nearest-neighbor convolution in GNNs. This work discusses a series of optimizations that have allowed scaling up the GFMs training to tens of thousands of GPUs on datasets consisting of hundreds of millions of graphs. Our GFMs use multi-task learning (MTL) to simultaneously learn graph-level and node-level properties of atomistic structures, such as energy and atomic forces. Using over 154 million atomistic structures for training, we illustrate the performance of our approach along with the lessons learned on two state-of-the-art United States Department of Energy (US-DOE) supercomputers, namely the Perlmutter petascale system at the National Energy Research Scientific Computing Center and the Frontier exascale system at Oak Ridge Leadership Computing Facility. The HydraGNN architecture enables the GFM to achieve near-linear strong scaling performance using more than 2,000 GPUs on Perlmutter and 16,000 GPUs on Frontier.
keywords:
Machine Learning, Atomistic Materials Modeling, Distributed Data Parallelism, Graph Foundation Models, Graph Neural Networks, Large-Scale Data Processing for Machine Learning
@@ 1 Introduction
Discovery of new materials with desired properties, accurate predictions of a material’s behavior throughout its entire lifespan is crucial to fundamental scientific progress in energy generation, transportation, electronics, and information technology [1]. Machine learning (ML) has shown great potential in accelerating the screening and pre-selection of materials for further experimental testing. In particular, deep learning (DL) models effectively captured relevant underlying relationships due to the arrangement of atoms of different constituents within various atomistic structures [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. DL models trained on data generated from experiments and/or first-principles calculations can predict the properties of interest to be used as new inputs. This inference takes only a fraction of the time it would take to run an experiment or a full first-principles calculation while still producing sufficiently accurate results. This drastic reduction in time to predict material properties using atomistic information results in a promising path towards accelerating material discovery and design [13, 14].
@@
However, generating vast volumes of experimental and/or first-principles training data is impractical even with sophisticated experimental facilities and powerful supercomputers. Recently, foundation models (FMs) have shown the promise to navigate around this challenge: once pre-trained over a large volume of available open-source data [15], an FM would provide a jump-start to refined models by fine-tuning on significantly smaller amounts of data for customized applications (also called downstream tasks). Reducing the number of simulations and/or experiments for generating domain-specific training data also drastically reduces the energy costs of developing domain-specific DL models.
@@
While state-of-the-art language-based FMs with a transformer architecture have reached promising results in several domains [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27], they fail to capture important topological aspects of the atomistic structures. Therefore, alternative DL architectures need to be considered for the development of trustworthy (interpreted as simultaneously accurate and highly confident) FMs for materials using atomistic scale information.
@@
Since atomistic material structures for a generic type of compound can be mapped onto a graph (where atoms can be treated as nodes and interatomic bonds as edges), graph foundation models (GFMs) [28, 29], which are FMs that operate on data structures as graphs, are the candidate of choice for these applications. Currently, GFMs proposed in the literature are developed by training graph neural network (GNN) architectures on a sufficiently large and comprehensive dataset for the domain of interest. While a few efforts have already been undertaken to develop GFMs for atomistic materials modeling applications [30, 31, 32, 33], the existing work is still at an incipient stage. Current efforts do not yet ensure that their proposed approach achieves trustworthiness. Moreover, serious concerns have been recently raised about the unaffordable amount of energy requested by proliferating AI centers due to computationally intensive tasks of training large FMs on large volumes of data.
@@
The work described in this manuscript developed scalable and trustworthy supervised GFMs for the simultaneous prediction of energies and atomic forces with careful attention to energy consumption. The GFMs have been constructed using HydraGNN [34, 35], a fully scalable GNN architecture developed at ORNL. Experiments were conducted on two large US-DOE supercomputers: the Perlmutter petascale machine at National Energy Research Scientific Computing Center (NERSC) and the Frontier exascale system at Oak Ridge Leadership Computing Facility (OLCF).
@@
The remaining content of the paper is organized as follows. We discuss the current state of the art and introduce HydraGNN in Section 2. In Section 3, we discuss our approach toward developing a scalable framework and list the different optimizations for scalable training. We discuss our use of large-scale HPO to develop a trustworthy GFM. Section 4 shows the performance of different components of this work: reading large data, scaling the training process, performing HPO at large scale, and measuring epistemic uncertainty using an ensemble of HPO trials with high predictive accuracy and low energy costs. We conclude our study and discuss future work in Section 5.
2 Current state-of-the-art
2.1 GNN training on open-source atomistic materials data
@@
To date, there have been a few approaches proposed in the literature to develop GFMs for atomistic materials modeling. In [30], the authors proposed a multi-modal approach where multiple encoding DL architectures are trained on different types of data representations and describing different physical quantities. The models are aligned to each other through a penalization term in the training loss function that forces latent vectors from each embedding space to coincide. However, the datasets used comprise only organic molecules, which cover only a relatively small set of natural elements on the periodic table.
@@
In [38], the authors propose a new kernel function for GNNs for modeling potential energy surfaces, which exploits chemical species information, and illustrate its efficacy in the context of transferable GFMs by pre-training the GFM on the Open Catalyst 2020 (OC2020) dataset [39] and illustrating its performance on a set of fine-tuning tasks. Since the GFM was pre-trained only on the OC2020 dataset, the applicability of this GFM is restricted to inorganic compounds.
While not explicitly presented by their developers as GFMs, there have been other models that cover broader sets of elements of the periodic table compared to the approaches mentioned in the previous paragraphs. In [31], the authors built a GNN model using MTL for simultaneous predictions of several material properties by training the GNN model on multiple datasets, including OC2020 [39] and Open Catalyst 2022 (OC2022) [40]. However, the approach considers only a single GNN architecture without performing HPO. Moreover, the set of parameters in the GNN model is relatively small, of the order of few millions of parameters, which limits the attainable accuracy on large volumes of data.
In [33], the authors studied the scaling behavior of 2D molecular GNNs under varied settings of depth, width, number of molecules, number of labels, the diversity in dataset, and architectural choice. The authors showed that supervised pre-training of large GNNs on molecular datasets provides a rich fingerprint embedding, which is useful for 38 downstream tasks. Even if this work very systematically studied the effect of GNN model size over the predictive performance in the pre-training and fine-tuning stage with many and diverse downstream tasks, the work has two important limitations: it only considers 2D graphs and it addressed only organic compounds.
Several UQ methods have been applied to GNNs [41], including Bayesian GNNs [42], prediction interval methods [43], and deep ensemble methods [44]. Bayesian methods are theoretically rigorous but challenging to scale to high-dimensional data. Prediction interval methods are cost-effective but often require tedious tuning of heuristic parameters.
Compared to the scientific contributions mentioned above, our work distinguishes itself by leveraging extreme scale supercomputing resources to ensure trustworthiness of the GFMs by performing (i) a systematic large scale HPO across a broad set of GNN architectures and (ii) a large scale ensemble learning (EL) for UQ, which realizes a compromise between cost and performance.
2.2 Scalability and GPU optimization for GNN training
@@
These are useful conclusions for optimization of GNN training on large graphs (i.e., with millions of nodes), but need to be re-evaluated for our datasets. Training on large graphs can be highly sensitive to the splitting scheme used to partition the graph into subgraphs and to distribute them among processors. For the atomistic materials modeling applications addressed in our work, the graph samples are small (with at most a few hundreds of nodes). For GNN convolutions in particular, convolution on a batch of samples will have a much more local, block diagonal structure. Therefore, throughput should be less sensitive to the choice of atomistic structures per batch.
@@
2.3 HydraGNN
The complexity of the physics and the scale at which atomistic structures must be studied in response to US-DOE needs in materials science makes it compelling to develop GNN capabilities that simultaneously satisfy several important algorithmic and computer science requirements. To effectively respond to the scientific needs of the US-DOE, a GNN architecture must provide: (1) capabilities to read and process data from multiple sources simultaneously, (2) flexibility to support diverse DOE-relevant scientific applications, (3) capabilities to scale the training on leadership class supercomputing facilities, (4) portability across heterogeneous computing environments, (5) continuous software maintenance by ensuring support and compatibility with upgraded software dependencies, and (6) maintained documentation to support new users across a broad set of international institutions.
While several GNN architectures have been made available as open-source tools to the scientific community in the last few years [51, 52, 53, 54], none of these tools satisfies all of the above requirements. Moreover, including missing capabilities on these well-established GNN libraries requires invasive and laborious modifications for software re-design. These challenges arising from existing GNN implementations motivated our effort in developing HydraGNN [34, 35]. In response to the US-DOE scientific needs, HydraGNN provides:
  • MTL capabilities to process multi-source, multi-fidelity data [55];
  • object-oriented programming capabilities to use different MPNN layers [56], which allows flexible switching between different message policies based on the scientific needs of the specific application at hand, as well treating the MPNN layer as a tunable categorical hyperparameter with HPO;
  • invariant and equivariant features that reduce computational redundancy and time-to-solution [57], therefore contributing to energy saving;
  • scalable input/output (I/O) data management techniques to efficiently scale the training of GNN models on millions of data samples using thousands of GPUs on supercomputing facilities [58]; and
  • portable capabilities that allow conveniently running the GNN training on diverse computing platforms with different hardware and software specifications.
@@ 2.3 HydraGNN
3 Our contribution
The work described in this manuscript is the first-of-its-kind large-scale training of GFMs for atomistic materials modeling using over 154 million atomistic structures as data samples and using over $91\%$ of the exascale supercomputer Frontier. We have used three key techniques for developing a scalable and trustworthy GFM: 1) scalable data management using a scientific data management library and an in-memory data store, 2) scalable HPO that uses asynchronous Bayesian optimization for efficiently managing computing resources, and 3) ensemble methods for UQ that allows model generalization and concurrently training multiple models. These three advancements collectively enhance the robustness, efficiency, and scalability of the GNN training process.
@@ 3.1 Data aggregation
Table 1: Overview of Datasets used for training HydraGNN
Using large datasets for GFM training can enhance generalizability and ensure resilience to data variance issues that typically arise during downstream tasks. To this end, we aggregated five open-source atomistic materials modeling datasets that are extremely diverse in terms of chemical composition, atomistic configurations, and number of atoms in the system. These datasets, as listed in Table 1, are: ANI1x, QM7x, OC2020, OC2022, and MPTrj.
  • ANI1x [63] consists of over 4,956,005 atomistic structures derived from up to 57 thousand distinct molecular configurations containing the C, H, N, and O chemical elements.
  • QM7x [64] is a comprehensive dataset of 42 physicochemical properties for approximately 4.2 million equilibrium and non-equilibrium structures of small organic molecules with up to seven non-hydrogen atoms from the C, N, O, S, Cl chemical elements.
@@
Each dataset is unique for the chemical compositions and the number of atoms in the atomistic structures of the compounds described. Figure 1 shows the distribution of the number of atoms and graph edges per atomistic structure for each dataset. For the MPTrj dataset, approximately half of the atomistic structures are relatively small in size. On the other hand, the OC2020 and OC2022 datasets consist of a more even distribution of atomistic structures with different sizes and edge counts, with the larger structures consisting of over 400 atoms and over 12,500 edges. In total, the data used for training, validating, and testing our GFM consisted of over 154 million atomistic structures that consume 5.3 Terabytes of storage space. These datasets were pre-processed using a scientific data management library into a common format for efficient storage and I/O, as discussed in Section 3.3.
@@
Figure 2 provides the heatmap that illustrates the frequency of occurrence of each element of the periodic table across the entire dataset that results from the aggregation of the datasets ANI1x, QM7-X, OC2020, OC2022, and MPTrj. We notice that there is an under-representation of the elements from the transition metal groups. This is partially due to the fact that first-principles calculations for materials including these elements (e.g., alloys) are more computationally expensive, and thus are available in smaller volumes with respect to first-principles data for other classes of materials.
@@
3.2 Data cleaning and pre-processing
Some of the atomistic structures were determined to have unrealistic values for atomic forces (on the order of 20,000 eV/angstrom). These probably corresponded to configurations visited at early stages of energy minimization during the generation of these datasets. To eliminate these outliers, we first applied a data cleaning operation in which we discarded all atomistic structures with an $L2$-norm (also known as spectral norm) of the force tensor above 100 eV/angstrom to ensure that these data samples did not affect the training of our GFMs. The number of data samples removed from each dataset by this filtering operation is reported in Table 2.